OpenStudy (xbeenchantedx):

how to do find the axis of symmetry for y=-1/4 - 4x?

OpenStudy (anonymous):

NO axis of symmetry

OpenStudy (xbeenchantedx):

oops i meant y=-1/4^2 - 4x

OpenStudy (anonymous):

Still no

OpenStudy (xbeenchantedx):

ugh i keep typing it wrong

OpenStudy (xbeenchantedx):

y=-1/4x^2 - 4x

OpenStudy (anonymous):

Diffrentiate and equate to zero

OpenStudy (xbeenchantedx):

can you show me how to do it?

OpenStudy (anonymous):

or it is x=(sum of roots)/2

OpenStudy (xbeenchantedx):

ok but for me i cant get the roots until i get the axis of symmetry. it turns into 4/(-2/4) for some reason

OpenStudy (anonymous):

sum of roots = -coefficient of x/coefficient of x^2 you don't need to find the roots

OpenStudy (xbeenchantedx):

so if i do do that it would be: sum of roots= -4/(-1/4) right?